OpenStudy (anonymous):

An object falls freely from rest near the surface of Earth, What is the speed of the object after having fallen a distance at 4,90 meters...how do i solve this?

OpenStudy (anonymous):

v^2 = vi^2 + 2as vi=0 in this case a= 9.81m/s^2 so V^2 = 2*9.81*490 V = 98.05 m/s